(y^2 +yz+ z^2)dx +(z^2 +zx+ x^2)dy + (x^2 +xy+ y^2)dz = 0
1
Expert's answer
2019-12-25T17:11:21-0500

Verify the Pfaffian Differential equation

"(y^2+yz+z^2)dx+(z^2+zx+x^2)dy+(x^2 +xy+y^2)dz=0"

is integrable and find its primitive.

The necessary and sufficient condition for iintegrability is


"X\cdot curlX=0,"

"X=(y^2+yz+z^2,z^2+zx+x^2,x^2 +xy+y^2)"

So that


"\nabla \times X=\begin{vmatrix}
 \mathbf i & \mathbf j & \mathbf k \\
 {\partial \over \partial x} & {\partial \over \partial y} & {\partial \over \partial z} \\
 y^2+yz+z^2 & z^2+zx+x^2 & x^2 +xy+y^2
\end{vmatrix}"

"=\mathbf i(x+2y-2z-x)- \mathbf j(2x+y-y-2z)+""+\mathbf k(z+2x-2y-z)""=2(y-z)\mathbf i+2(z-x) \mathbf j+2(x-y) \mathbf k"

"X\cdot curlX=2(y^2+yz+z^2)(y-z)+""+2(z^2+zx+x^2)(z-x)+2(x^2 +xy+y^2)(x-y)=""=2(y^3-y^2z+y^2z-yz^2+yz^2-z^3)+""+2(z^3-xz^2+xz^2-x^2z+x^2z-x^3)+""=2(x^3-x^2y+x^2y-xy^2+xy^2-y^3)=0"

Thus the given equation is integrable.


"P=y^2+yz+z^2,""Q=z^2+zx+x^2,""R=x^2 +xy+y^2."

The auxiliary equations are


"{dx \over {\partial Q\over \partial z}-{\partial R \over \partial y}}={dy \over {\partial R\over \partial x}-{\partial P \over \partial z}}={dz \over {\partial P\over \partial y}-{\partial Q \over \partial x}}"

"{dx \over 2(z-y)}={dy \over 2(x-z)}={dz \over 2(y-x)}"

"{dx \over z-y}={dy \over x-z}={dz \over y-x}"


"dx+dy+dz=0=>x+y+z=c_1=u"

"(y+z)dx+(x+z)dy+(x+y)dz=0=>xy+yz+zx=c_2=v"

Formulate the equation "Adu+Bdv=0" and compare with the given equation.

"Adx+Ady+Adz+Bxdy+Bydx+Bzdy+Bydz+Bzdx+Bxdz=0"


"A+By+Bz=y^2+yz+z^2""A+Bx+Bz=z^2+xz+x^2""A+By+Bx=x^2+xy+y^2"

"B=x+y+z""A=-xy-yz-xz"

Then


"A=-v""B=u"

"Adu+Bdv=0""-vdu+udv=0"

"{dv \over v}={du \over u}"

"\ln v=\ln u+\ln c""v=cu"

The required solution is


"xy+yz+zx=c(x+y+z)"